OpenStudy (anonymous):

What is the value of b in the equation 3b + 2(b − 1) = 8? A. 2 B. 5 C. 8 D.10 HELPPPP

OpenStudy (madgirlwithabluebox):

i believe it is a

OpenStudy (anonymous):

Thank youuu

OpenStudy (madgirlwithabluebox):

3(2)+2(2-1)=8 6+2(1)=8 6+2=8 8=8

OpenStudy (madgirlwithabluebox):

Incase you needed my work

OpenStudy (anonymous):

it is a or c ? amd thanks so much

OpenStudy (anonymous):

@Madgirlwithabluebox

OpenStudy (triciaal):

the correct way to solve first expand the parenthesis group like terms divide by the coefficient of the variable

OpenStudy (madgirlwithabluebox):

its a. because you plug in 2 for b

OpenStudy (triciaal):

@Madgirlwithabluebox yes a is correct b = 2 that's from substituting the given choices

OpenStudy (anonymous):

What is the value of y in the equation 3y − 8 = 22? A. 7.3 B.8 C. 10 D. 14

OpenStudy (triciaal):